Circular No. 6740

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

USNO 1425.09823278 T. Vanmunster, Landen, Belgium, reports that unfiltered CCD obser-

vations during a 7.8-hr run on Sept. 10–11 with a 0.25-m telescope reveal superhumps in this system (cf. IAUC 6731), with a mean semiamplitude of 0.05 mag, hence classifying it as an SU UMa-type dwarf nova. A period analysis using the phase dispersion minimization method yields a superhump period of 0.0561 ± 0.0004 day.

Visual magnitude estimate by K. Sárneczky, Ráktanya, Hungary: Sept. 5.852 UT, 13.8.

V518 PERSEI A. Guarnieri, C. Bartolini, N. Masetti, and A. Piccioni, University

of Bologna, communicate: “The analysis of one CCD image of V518 Per obtained with the Bologna University 1.5-m telescope (+ BFOSC) in the Cousins R band on Jan. 16.85 UT, during an x-ray mini-outburst (IAUC 6605), shows the object at R_{c} = 20.7 ± 0.1. This is at least 0.4 mag brighter than the quiescent R_{c} value reported by Chevalier and Ilovaisky (1996, A.Ap. 312, 105) and indicates that the periodic x-ray activity of the system (IAUC 6738) is also present in the optical.”

COMET C/1997 J1 (MUELLER) Ephemeris extension to IAUC 6672 (orbital elements MPC 30428):

1997 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Aug. 30 9^{h}28.^{m}62 +40^{o}00.^{′}5 3.426 2.652 34.^{o}2 12.^{o}4 13.4
Sept. 9 9 30.21 +38 53.0 3.395 2.708  40.4 14.0 13.5
19 9 30.78 +37 57.8 3.339 2.766  47.8 15.6 13.5
29 9 30.01 +37 14.5 3.260 2.827  56.1 17.1 13.6
Oct. 9 9 27.47 +36 43.4 3.162 2.890  65.3 18.3 13.6
19 9 22.68 +36 23.9 3.049 2.955  75.2 19.0 13.6
29 9 15.11 +36 15.0 2.927 3.022  85.9 19.1 13.6
Nov. 8 9 04.13 +36 14.6 2.802 3.091  97.5 18.5 13.6
18 8 49.21 +36 18.1 2.685 3.161 109.9 17.1 13.6
28 8 30.02 +36 18.6 2.584 3.233 123.3 14.8 13.7
Dec. 8 8 06.76 +36 06.5 2.513 3.306 137.4 11.6 13.7
18 7 40.49 +35 31.7 2.480 3.380 151.9  7.9 13.8
28 7 13.06 +34 28.5 2.495 3.454 165.0  4.2 13.9
1997 September 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green